OpenStudy (anonymous):

Solve by the substitution method. What is the solution? Type as an ordered pair. 2m + n = 10 m - 4n = 14

OpenStudy (anonymous):

n=`0-2m substitute in the second equation: m-4(10-2m)=14 m-40+8m=14 9m=54 m=6 n=10-2m=10-12=-2 (m;n)=(6;-2)

OpenStudy (anonymous):

So the odered pair is (6, -2)?